In R

4

let L

1

be the subspace spaned by the vectors a=(1;2;5;1), b=(4;3;3;0), c=(7;4;1;−1), and let L

2

be the subspace spaned by the vectors d=(1;1;1;1), f=(−1;0;3;−1) and g=(5;2;−1;−3). Find the dimension of L

1

∩L

2

.


1
Expert's answer
2021-10-08T15:28:02-0400

"dim (L_1\cap L_2)=dimL_1+dimL_2-dim(L_1\cup L_2)." We can see that "dimL_1=2" , because the rank of the matrix composed of the coordinates of the vectors "a,b,c" equals to 2:

"rank \begin{pmatrix} 1 & 2 &5& 1\\
 4 & 3 &3&0\\
 7& 4& 1& -1\\ \end{pmatrix}=" "rank \begin{pmatrix}
 1 & 2 &5& 1\\
 0 & -5 &-17&-4\\
 0& -10& -34& -8 
\end{pmatrix}=" "rank \begin{pmatrix}
 1 & 2 &5& 1\\
 0 & -5 &-17&-4
\end{pmatrix}=2" .

We can see that "dimL_2=3" , because

"rank \begin{pmatrix}
 1 & 1 &1& 1\\
 -1 & 0 &3&-1\\
 5& 2& -1& -3 
\end{pmatrix}=" "rank \begin{pmatrix}
 1 & 1 &1& 1\\
 0 & 1 &4&0\\
 0& -3& -6& -8 
\end{pmatrix}=" "rank \begin{pmatrix}
 1 & 1 &1& 1\\
 0 & 1 &4&0\\
 0& 0& 6& -8 
\end{pmatrix}=3"

If we want find the "dim(L_1\cup L_2)" , we can find the rank of a matrix composed of vectors "a,b,c,d,f,g" :

"rank \begin{pmatrix}
 1 & 2 &5& 1\\
 4 & 3 &3&0\\
 7& 4& 1& -1\\ 1 & 1 &1& 1\\
 -1 & 0 &3&-1\\
 5& 2& -1& -3 
\end{pmatrix}=" "rank \begin{pmatrix}
 1 & 2 &5& 1\\
 0 & -5 &-17&-4\\
 0& -10& -34& -8 \\ 0 & -1 &-4& 0\\
 0 &2 &8&0\\
 0& -8& -26& -8 
\end{pmatrix}="  "rank \begin{pmatrix}
 1 & 2 &5& 1\\ 0 & -1 &-4& 0\\
 0 & -5 &-17&-4\\
 0& -8& -26& -8 
\end{pmatrix}=" "rank \begin{pmatrix}
 1 & 2 &5& 1\\ 0 & -1 &-4& 0\\
 0 & 5 &17&4\\
 0& 4& 13& 4 
\end{pmatrix}=" "rank \begin{pmatrix}
 1 & 2 &5& 1\\ 0 & -1 &-4& 0\\
 0 & 0 &-3&4
\end{pmatrix}=3"

So "dim(L_1\cup L_2)=3". It follows that  "dim (L_1\cap L_2)=2+3-3=2" .